OpenStudy (unicwaan):

Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (freckles):

compare the real parts of both sides and compare the imaginary parts of both sides a+bi a is the real part b is the imaginary part so what is the real part of -3+yi and the real part of x+6i?

OpenStudy (unicwaan):

-3 and x is the real part?

OpenStudy (freckles):

so x=-3

OpenStudy (freckles):

so now compare the coefficients of the imaginary unit of i of both sides that is what is the imaginary part of -3+yi and x+6i?

OpenStudy (unicwaan):

yi and 6i

OpenStudy (freckles):

so what must y be so that yi=6i

OpenStudy (unicwaan):

y = (6i)/i? can I simiplfy that?

OpenStudy (unicwaan):

simplify*

OpenStudy (freckles):

well looks like you are trying to solve for y even though there is really nothing to solve but I suppose you can do this longer way it isn't too much longer anyways but yeah what is i/i ?

OpenStudy (unicwaan):

1, so the i's cancel out and I'm just left 6?

OpenStudy (freckles):

yep but you could have skipped all of that yi=6i is only satisfied when y is 6

OpenStudy (unicwaan):

okay, thank you so much!!

OpenStudy (freckles):

no problem in general if you have (Let x,y,a,b be real numbers) x+yi=a+bi then x=a and y=b

OpenStudy (unicwaan):

So for future refernece I just have to separate the real and imaginary parts and solve for the variable, that makes sense, thanks!

OpenStudy (freckles):

yep yep example: 3+yi=x-9i+2i anyways first write this as 3+yi=x-7i so x=3 and y=-7

OpenStudy (unicwaan):

thank you :)

OpenStudy (freckles):

np :)